### v3

Abstract
Python serves as an open-source and cost-effective alternative to the MATLAB programming language. This paper introduces a concise topology optimization Python code, named “PyHexTop,” primarily intended for educational purposes. Code employs hexagonal elements to parameterize design domains as such elements provide checkerboard-free optimized design naturally. PyHexTop is developed based on the “HoneyTop90” MATLAB code [1] and uses the NumPy and SciPy libraries. Code is straightforward and easily comprehensible, proving a helpful tool that can help people new in the topology optimization field to learn and explore. PyHexTop is specifically tailored to address compliance minimization with specified volume constraints. The paper provides a detailed explanation of the code for solving the Messerschmitt-Bolkow-Blohm beam and extensions to solve problems different problems. The code is publicly shared at: https://github.com/PrabhatIn/PyHexTop.
Keywords:
Topology optimization, Honeycomb tessellation, Python, Compliance minimization
1 Introduction
Topology optimization (TO) concerns the best locations to place cavities in a structure design domain by extremising a desired objective with a given set of constraints. This field has undergone extensive research over the past three decades, resulting in numerous research articles that propose various TO algorithms for implementing methods like Solid Isotropic Material with Penalization (SIMP) [2], Level-set method [3], evolutionary structural optimization method [4], and feature-based [5, 6] approach, etc. Providing educational code in TO is a positive trend initiated by Sigmund [7] with the 99-line MATLAB code. Making such codes publicly available helps the field’s growth to different regimes and applications, wherein newcomers find it easy to learn, explore, and extend. Python has emerged as a cost-effective, open-source alternative to the MATLAB programming language and is gaining popularity. Only a few published TO codes using Python interfaces exist [8, 9]. This work introduces a standalone Python code called “PyHexTop” for educational purposes, explicitly targeting newcomers and students. The code uses hexagonal elements to parameterize the design domain. These elements are generated by method outlined in HoneyTop90 MATLAB code [1]. Such elements provide edge connections between two adjacent elements, eliminating checkerboard patterns from the optimized designs [1, 6].
Python is a dynamic and versatile programming language with great utility in various domains, including TO. Its powerful libraries and frameworks, like NumPy and SciPy, are best suited for scientific computation, data analysis, and numerical optimization. Thus, Python can be an excellent choice for implementing the TO technique. The code presented in this paper offers a user-friendly platform that encourages further development and flexibility to implement alternate optimization methods with different underlying physics.
(a) MBB beam
(b) Numbering scheme
(c) Multi-load cases
(d) $R_{0}$: Void region, $R_{1}$: Solid region
Figure 1: (1(a)) A symmetric MBB beam design domain. (1(b)) A schematic for node (blue texts) and element (red texts) numbering scheme [1]. (1(c)) Multiple load cases domain. Texts in circles indicate load cases. (1(d)) Domain with passive regions
2 Problem Formulation
The MBB beam design problem is considered to demonstrate and explain the presented Python code. Fig. 1(a) depicts the symmetric half-design domain of the beam with the pertinent boundary conditions and external load F.
[…]
$\mathbf{k}_{j}=E(\rho_{j})\mathbf{k}_{0}=\left(E_{\text{min}}+\rho_{j}^{p}\left(E_{0}-E_{\text{min}}\right)\right)\mathbf{k}_{0},$ (1)
$E(\rho_{j})$, $E_{0}$ and $E_{\text{min}}$ indicate Young’s moduli of element $j$, $\rho_{j}=1$ (solid), $\rho_{j}=0$ (void), respectively. Material contrast i.e. $\frac{E_{\text{min}}}{E_{0}}=10^{-9}$ is taken [1]. $\mathbf{k}_{0}$ is the element stiffness matrix with $E(\rho_{j})=1$. $p$, penalty parameter, is set to 3. We solve the following optimization problem:
$\begin{rcases}\begin{split}&{\min_{\bm{\rho}}}\quad C({\bm{\rho}})=\mathbf{u}^{\top}\mathbf{K}(\bm{\rho})\mathbf{u}=\sum_{j=1}^{\mathtt{Nelem}}\mathbf{u}_{j}^{\top}\mathbf{k}_{j}(\rho_{j})\mathbf{u}_{j},\quad\bm{0}\leq\bm{\rho}\leq\bm{1},\,\text{with}\\ &(i)\,\,\bm{\lambda}:\,\,\mathbf{K}\mathbf{u}-\mathbf{F}=\mathbf{0},\,(ii)\,\,\Lambda:\,\,\text{g}=\frac{V(\bm{\rho})}{\mathtt{Nelem}\,V^{*}_{f}}-1=\frac{\sum_{j=1}^{\mathtt{Nelem}}v_{j}\rho_{j}}{\mathtt{Nelem}\,V^{*}_{f}}-1\leq 0,\end{split}\end{rcases}$ (2)
where $C({\bm{\rho}})$ is compliance of the structure. $\mathbf{u}$ and $\mathbf{F}$ indicate the global displacement and force vectors. $\mathbf{u}_{j}$ is displacement vector of element $j$. $V$ and $V^{*}_{f}*\texttt{Nelem}$ are the optimized designs’ total material and desired volumes. $\bm{\rho}$, design vector, contains all $\rho_{j}$. $\bm{\lambda}$ (vector) and $\Lambda$ (scalar) are the Lagrange multipliers (Eq. 2). $\bm{\lambda=-2\mathbf{u}}$ [1]. Sensitivities of $C$ wrt. $\rho_{j}$ is $\frac{\partial C}{\partial\rho_{j}}=-\mathbf{u}_{j}^{\top}\frac{\partial\mathbf{k}_{j}}{\partial\rho_{j}}\mathbf{u}_{j}=-p(E_{0}-E_{\text{min}})\rho_{j}^{p-1}\mathbf{u}_{j}^{\top}\mathbf{k}_{0}\mathbf{u}_{j}$ [1]. Likewise, one finds derivatives of the volume constraint $\frac{\partial\text{g}}{\partial\rho_{j}}=\frac{v_{j}}{\mathtt{Nelem}\,V^{*}_{f}}=\frac{1}{\mathtt{Nelem}\,V^{*}_{f}}$, wherein $v_{j}=1$ is assumed. The standard optimality criteria [7] is used to solve the optimization problem (Eq. 2). We use sensitivity [7], density [10], and null filterings [1] to solve the problems presented.
3 Python Implementation Detail
This section presents the Python implementation of PyHexTop in detail. We use the method outlined in [1] to generate the element connectivity and nodal coordinates matrices. The MBB beam is optimized to demonstrate the said code.
3.1 Element Connectivity and Nodal Coordinates matrices (lines 15-28)
Lines 3-8 of PyHexTop are needed to install all the required libraries. Let HNex and HNey be the number of hexagonal elements in $x$ and $y$ directions, respectively. Each element consists of six nodes. Each node contains two degrees of freedom (DOFs). The Numpy array NstartVs (line 15) assigns node numbers of elements row by row, starting from the bottom (See 1(b)). DOFstartVs (line 16) stores the DOFs for the symmetrical half quadrilaterals of each hexagonal element, and the order of storage follows the numbering of the nodes, with each node’s DOFs arranged sequentially. This array, modified for DOF organization, is horizontally replicated eight times and then vertically replicated $2\times\texttt{HNey}\times\texttt{HNex}$ times. The resulting matrix is named NodeDOFs (line 17). The redundant DOFs related to symmetry are removed using the setdiff Numpy function, and the resulting matrix is stored in ActualDOFs (line 18). The final connectivity matrix HoneyDOFs (line 19) contains organized DOFs relevant to the honeycomb structure, with the specified columns from ActualDOFs. Readers are recommended to refer to [1] for more detail.
The $y$ coordinates of the nodes for a honeycomb tessellation are determined on lines 20-22 and stored in array Ncyf. Similarly, the $x$ coordinates are stored on line 24. The matrix HoneyNCO (line 25) records the final nodal coordinates, wherein $x$ and $y$ coordinates are stored in the first and second columns, respectively. When HNey is an even number, hanging nodes are found and removed [1]. HoneyDOFs and HoneyNCO are updated on lines 27 and 28, respectively.
3.2 Finite Element Analysis (lines 31-45 and 86-89)
The total number of elements Nelem and nodes Nnode for the current mesh grid are calculated on line 29. Fixed DOFs are determined and stored in array fixeddofs (line 37). Total DOFs are listed in alldofs (line 38). The array freedofs (line 39) identifies the available DOFs by excluding the fixed DOFs from the set of alldofs by using the Numpy setdiff1d function.
A sparse force array $\mathbf{F}$ is initialized on line 34 using SciPy function csc_matrix. The initialization involves assigning a non-zero force value from the data array at the indices specified by row and col vectors. The csc_matrix allows for efficient sparse matrix initialization. The displacement vector U is initialized on line 36. On lines 41-44, the rows and column indices of the matrix HoneyDOFs are recorded in vectors iK and jK respectively, using the kron Numpy function. The reshaping and type conversion ensures the indices are appropriately formatted for further use. E0 (line 12) and Emin (line 13) are stored. Elemental stiffness matrix KE (lines 45-56) is evaluated using Wachpress shape functions with plane stress assumptions [1]. The global stiffness matrix K is determined (line 87). The finite element analysis is performed on lines 86-89 using a sparse linear solver spsolve and U is updated on line 89 with the freed DOFs.
3.3 Filtering, Optimization and Results Printing (lines 58-81 and and 91-122 )
Three filtering options are provided: sensitivity filtering $\mathtt{(ft=1)}$, density filtering $\mathtt{(ft=1)}$, and null (no) filtering $\mathtt{(ft=0)}$. The latter is included to illustrate the behavior of hexagonal elements for checkerboard patterns and point connections. We calculate the centroid coordinates of FEs in lines 58-60 (Cxx and Cyy). These vectors are combined in the matrix ct on line 61.
With rfill, a loop is run to identify neighboring elements within the filter radius, and their indices are recorded in I (line 68). Concurrently, a corresponding index array J is created (line 69). The $x,\,y,$ and distance information of neighboring elements, along with their respective indices, are consolidated into arrays newx, newy, and newz, respectively. Arrays are combined to form the matrix DD (lines 71-74) containing the $x$ and $y$ coordinates and distance of neighboring FEs. The filtering matrix HHs is evaluated on lines 76-77 using spdiags and coo_matrix SciPy that create a sparse diagonal matrix with indices based on the distances in DD.
volfrac is used for the initial guess of TO (line 79). $\rho$ is denoted by x. xPhys represents the physical material density vector (line 79). With either $\mathtt{ft=0}$ or $\mathtt{ft=1}$, xPhys is equal to x, whereas with $\mathtt{ft=2}$, it is represented via the filtered density on line 113. loop, change, maxiter, dv and move are defined on line 80. The element compliance ce for each FE is calculated on line 91. c is calculated using material properties and xPhys. The vector dc calculates the sensitivity of c wrt. xPhys, and is later updated as per different ft.
The optimality criteria for TO is implemented on lines 101-114. xOpt stores the original design variable. xUpp and xLow define the upper and lower bounds of x, respectively. OcC computes the normalized change in x. inL gives an interval for the optimization search. Vector dv records volume constraint sensitivities. The loop iteratively updates design variables to achieve the desired volume fraction. The convergence and termination conditions are monitored by the vector change. A scatter plot is created to visualize the result using xPhys for color intensity.
(a) C = 298.92
(b) C = 293.66
(c) C = 282.47
(d) C = 332.27
(e) C = 335.65
(f) C = 320.53
(g) C = 365.54
(h) C = 373.65
(i) C = 356.31
Figure 2: Optimized MBB beams. The results in the first (a-c), second (d-f), and third (g-i) rows are displayed for ft=0, ft=1, and ft=2 respectively.
3.4 MBB beam optimized results
PyHexTop code is called as:
PyHexTop(HNex, HNey, rfill, volfrac, penal, ft)
The selected mesh sizes (HNex $\times$ HNey) are $60\times 20$, $120\times 40$ and $300\times 100$ with the filter radius rfill, $2.4\sqrt{3}$, $4.8\sqrt{3}$, and $12\sqrt{3}$, respectively, i.e., 0.04 times the maximum length. volfrac is set to 0.5, and penal is set to 3.
The results in Fig. 2 show that designs with $\mathtt{ft=0}$ (1st row) are the best performing and without checkerboard patterns, albeit being mesh-dependent and featuring thin members. To address this, either filtering or length scale constraints [11] can be used. The designs in 2nd and 3rd rows are achieved with $\mathtt{ft=1}$ and $\mathtt{ft=2}$, respectively. These designs eliminate checkerboard patterns and remain mesh-independent, maintaining consistent topology regardless of the mesh sizes.
(a) C = 92.27
(b) C = 252.92
(c) C = 156.35
(d) C = 160.55
Figure 3: Optimized results: multi-load cases (3(a)) 2-load cases (3(b)) 4-load cases; Passive regions (3(c)) $\mathtt{ft=1}$ (3(d)) $\mathtt{ft=2}$
4 Simple extensions
This section provides extensions of the Python code for different problems.
4.1 Multiple loads
PyHexTop code is modified to solve a cantilever beam design with different multi-load cases (Fig. 1(c)). $F$ is adjusted to a column vector per the number of load cases. Accordingly, $U$ is calculated and stored in a two or four-column format. $C=\sum_{k=1}^{2}\mathbf{U}_{k}^{\top}\mathbf{K}\mathbf{U}_{k}$ where $\bm{U_{k}}$ is the displacement for the $k^{th}$ load case. For four load cases, the following modifications are made:
row = np.array([(2*HNex+1)*2 - 1, 2*Nnode-1, 2*(HNex+1) -1, 2*(HNex + HNey + 2*HNex*HNey) - 1]);col = np.array([0,1,2,3]); data = np.array([-1,1,2,-2]);F = csc_matrix((data, (row, col)), shape=(2*Nnode, 4))
The U and fixeddofs matrices are adjusted as
U = np.zeros((2*Nnode, 4));fixeddofs = np.append(2 * (np.arange(1, (2*HNex+1)*HNey+2, 2*HNex+1))-1, 2*(np.arange(1,(2*HNex+1)*HNey+2, 2*HNex+1)))
The following changes are done to determine objective and sensitivity:
c=0;dc = 0
for i in range(0,(F.shape[1])):
Ui = U[:,i];ce = np.sum((Ui[HoneyDOFs-1].dot(KE))*(Ui[HoneyDOFs-1]),1)
c = c+sum((Emin+xPhys**penal*(E0-Emin))* (ce.reshape((-1, 1)))).sum()
dc = dc - penal*(E0 - Emin) * xPhys ** (penal-1)*(ce.reshape((-1, 1)))
The code is called as below.
PyHexTop(120, 120,4 * np.sqrt(3), 0.4, 3, 1)
The optimized designs for 2-load and 3-load cases are depicted in Figs. 3(a) and 3(b).
4.2 Passive design domains
A problem [1] with passive (non-design) regions is shown in Fig. 1(d). These modifications are made to code for the problem. The force and fixed DOFs vectors are changed to
row = np.array([2*(2*HNex+1)-1]);col = np.array([0]);data = np.array([-1])
F = csc_matrix((data, (row, col)), shape=(2*Nnode, 1))
fixeddofs = np.append(2 * (np.arange(1, (2*HNex+1)*HNey+2, 2*HNex+1))-1, 2*(np.arange(1, (2*HNex+1)*HNey+2, 2*HNex+1)))
Array RSoVo, is initialised to zeros(Nelem,1) after line 65 to store information related to the regions $\mathbf{R_{1}}$ (solid) and $\mathbf{R_{0}}$ (void). RSoVo is then modified as follows (line 74-77)
if (np.sqrt((ct[j,0]-np.max(ct[:,0])/3)**2+(ct[j,1]-np.max(ct[:,1])/2)**2)<np.max(ct[:,1])/3):
RSoVo[j] = -1 #non-design void region
elif (ct[j,0]>0.7*np.max(ct[:,0]) and ct[j,0]<0.9* np.max(ct[:,0]) and ct[j,1]>0.1*np.max(ct[:,1]) and ct[j,1]<0.3* np.max(ct[:,1])):
RSoVo[j] = 1 #non-design solid region
FEs with $\texttt{RSoVo}=-1$ or $\texttt{RSoVo}=1$ are passive elements. Vector act is included after line 77 to store active FEs. NVe and NSe are arrays for the $\mathbf{R_{0}}$ and $\mathbf{R_{1}}$ regions respectively.
NVe,NSe = np.where(RSoVo == -1), np.where(RSoVo == 1)
act=np.setdiff1d(np.arange(1,Nelem + 1).reshape(-1, 1),np.union1d(NVe,NSe))
x [act-1] = (volfrac * (Nelem - len(NSe)) - len(NVe))/len(act)
x[act-1] contains initialized design variables per act. The sensitivities of the objective and constraints wrt. the passive variables are zero. Only active sensitivities are taken on line 113 for performing optimization. The following code is added after density filtering on line 103 for accounting non-design regions.
xPhys[RSoVo == 1] = 1; xPhys[RSoVo == -1]= 0
The code is called as below and optimized designs for ft=1 and ft=2 are displayed in Figs. 3(c) and 3(d), respectively.
PyHexTop(200,100, 5.6* np.sqrt(3), 0.4, 3, ft)
5 Closure
This paper provides a Python code, named PyHexTop, for TO with hexagonal elements. Such elements provide edge connectivity between two juxtaposed elements, thus helping achieve checkerboard-free optimized designs. The code aims to facilitate newcomers to learn, explore, and develop codes for various applications. Efficient use of Python’s NumPy and SciPy libraries allows to generate nodal coordinates and element connectivity matrix per [1]. Stiffness matrix of elements is calculated using Wachpress shape functions. PyHexTop employs the optimality criteria approach to solve compliance minimization problems.
PyHexTop and its extensions are discussed, offering a comprehensive understating of the code’s implementation. Moreover, simplicity and flexibility of PyHexTop open up several possibilities for future research and developments, e.g., problems with stress/buckling constraints and design-dependent loads [12, 13].